Answer the following worded problems below. Show your solutions.


A 17.3L of gas enclosed in a container having a standard pressure which is 760mmHg. If the pressure is lowered to 240mmHg, what will be the volume of the gas?


A 63L of HCl is heat-cured at 102.6 degrees Celsius, by adding another 63.44L of the same substance, the temperature lowers. What will be the temperature after adding the additional volume of HCl?


A series of changes in pressure in an oxygen tank, starting at 12kPa to 42kPa and to 56kPa. What will be the total pressure if the last state of oxygen inside the tank is thrice the initial pressure?


6L of water reached its boiling point at 46kPa. If the pressure increases at 99kPa and the volume lowers at 2.8L, what is the change in temperature?



1
Expert's answer
2022-01-21T05:41:02-0500

1) P1V1 = P2V2 ; V2 = P1V1 / P2; V2 = (17.3L ×  760mmHg) / 240mmHg = 54.7L


2) V1T2 = V2T1 ; T2 = V2T1 / V1; T2 = (63.44L × 102.6°) / 63L = 103.3°


3) Ptotal = P1 + P2 + P3 + P4

P4 = 3P1 = 3 × 12kPa = 36kPa

Ptotal = 12kPa + 42kPa + 56kPa + 36kPa = 146kPa


4) PT/V = const

P1T1/V1 = P2T2/V2; (46/6)T1 = (99/2.8)T2; 7.6T1 = 35.35T2

ΔT = Т2 / T1 = 35.35 / 7.6 = 4.6 (K)